# High-Level Synthetic Data Generation with Data Set Archetypes

### v3

@@ Abstract
Cluster analysis relies on effective benchmarks for evaluating and comparing different algorithms. Simulation studies on synthetic data are popular because important features of the data sets, such as the overlap between clusters, or the variation in cluster shapes, can be effectively varied. Unfortunately, curating evaluation scenarios is often laborious, as practitioners must find lower-level geometric parameters (like cluster covariance matrices) to match a higher-level scenario description like “clusters with very different shapes.” To make benchmarks more convenient and informative, we propose synthetic data generation based on data set archetypes. In this paradigm, the user describes an evaluation scenario in a high-level manner, and the software automatically generates data sets with the desired characteristics. Combining such data set archetypes with large language models (LLMs), it is possible to set up benchmarks purely from verbal descriptions of the evaluation scenarios. We provide an open-source Python package, repliclust, that implements this workflow. A demo of data generation from verbal inputs is available at demo.repliclust.org.
@@ 1 Introduction
The goal of clustering is to separate data points into groups such that points within a group a more similar to each other than to those outside the group (Cormack, (1971)). In practice, it is often not clear what constitutes a cluster (Hennig, (2015)). As a result, many practitioners evaluate cluster analysis algorithms on synthetic data (Milligan, (1980), Milligan et al., (1983), Tibshirani et al., (2001), Steinley and Brusco, (2008), Steinley and Brusco, (2011), Van Mechelen et al., (2023)).
Synthetic data is valuable for two reasons. First, it clearly stipulates which data points belong to which cluster, allowing objective evaluation. Second, it allows independently manipulating different aspects of the data (such as the overlap between clusters or the variability of cluster shapes), which is critical for drawing scientific conclusions about the relative merits of different cluster analysis techniques (Milligan, (1996)).
Unfortunately, setting up benchmarks with synthetic data can be laborious. The process typically involves creating data sets for a number of different scenarios. For example, on benchmarks with convex clusters drawn from probabilistic mixture models, the scenarios may involve “clusters of very different shapes and sizes”, “highly overlapping oblong clusters”, “high-dimensional spherical clusters”, etc. (Tibshirani et al., (2001)).
Existing data generators do not cater directly to such high-level scenarios. Instead, the user must carefully tune simulation parameters to arrive at the desired scenarios (Steinley and Henson, (2005), Schubert and Zimek, (2019), Iglesias et al., (2019)). While some generators make it easy to control the overlaps between clusters, such high-level control typically does not extend to other aspects like the desired diversity of cluster shapes and sizes (Qiu and Joe, (2006), Shand et al., (2019)).
@@ 2 Generating Data from High-Level Archetypes
We implement archetype-based generation by summarizing the overall geometry of probabilistic mixture models in terms of a few high-level parameters, as listed in Table 1. To create an archetype, users can either specify these parameters directly or verbally describe the archetype in English. To map an English description to a precise parameter setting, we use few-shot prompting of a large language model (Brown et al., (2020), OpenAI et al., (2024)).
@@ 3.1 Defining Clusters and Mixture Models
[…]
@@ 3.2.1 Measuring Pairwise Overlap
Steinley and Henson, (2005) quantify cluster overlap by computing the full distributional overlap in $p$ dimensions. We prefer our one-dimensional notion in terms of minimax classification error, since high-dimensional Euclidean geometry exhibits a number of counterintuitive phenomena. Specifically, as $p\rightarrow\infty$, the majority of a sphere’s volume becomes concentrated in a thin shell from its surface, and so $p$-dimensional overlap goes to zero unless the marginal overlaps in each dimension approach 100% (Blum et al., (2020), Steinley and Henson, (2005)).
Figure 3: Cluster overlap based on the misclassification rate of the best linear classifier, in a minimax sense. The left panel shows the densities of two normal distributions in 1D. A minimax classification rule implies that the conditional misclassification rates for the blue and red distributions are equal. The black dashed line is the resulting decision boundary. The gray shaded areas on each side of this line represent classification mistakes. When viewed as one-dimensional clusters, the overlap is the total probability mass in the gray areas. The right panel shows two multivariate normal clusters in 2D. The black dashed line is the decision boundary based on the minimax classification axis $\boldsymbol{a}^{*}$. The cluster overlap is $\alpha=14.7\%$.
@@
For a pair of multivariate normal clusters with means $\boldsymbol{\mu}_{1}\neq\boldsymbol{\mu}_{2}$ and the same covariance matrix $\boldsymbol{\Sigma}$, the axis $\boldsymbol{a}_{\text{\tiny{LDA}}}=\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$ minimizes the overall misclassification rate, assuming equal class probabilities (see Hastie et al., (2009)). Unfortunately, the result does not hold for unequal covariance matrices $\boldsymbol{\Sigma}_{1}\neq\boldsymbol{\Sigma}_{2}$. Thus, we propose the approximation $\boldsymbol{a}_{\text{\tiny{LDA}}}=(\frac{\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}}{2})^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$. Figure 4 verifies that this LDA-based approximation closely matches the exact overlap, as compared with a cruder “center-to-center" (C2C) approximation that uses the difference between the cluster centers as the classification axis (i.e., $\mathbf{a}_{\tiny{C2C}}=\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1}$).
Figure 4: Quality of approximating cluster overlap using our LDA and the cruder center-to-center (C2C) approximations. Each data point corresponds to a pair of multivariate normal clusters with the pairwise overlap shown. The full simulation is based on 900 pairs generated from a variety of data set archetypes with different cluster shapes and numbers of dimensions. The dashed lines estimate the conditional means. The C2C and LDA approximations are both highly correlated with exact cluster overlap, with Pearson correlations $r$ close to 1 (left). However, the C2C method incurs significant relative error, while the LDA approximation typically comes within 10% of the exact overlap (right).
@@ 3.4 Making Cluster Shapes More Irregular
In many application domains, ellipsoidal center-based clusters are good models for the data. This is often the case when clusters can be characterized by a “typical” element. For example, in single-cell RNA sequencing, Chen et al., (2020) represent the differences in transcriptional activity across cell types using Gaussian mixture models. In this case, the cluster centers correspond to each cell type’s typical gene expression profile. Similarly, in deep learning, a large language model’s hidden states often form compact clusters that can effectively classify text (Howard and Ruder, (2018)), assess the truthfulness of a response (Azaria and Mitchell, (2023)), and help detect out-of-distribution inputs (Ren et al., (2023)).
However, convex clusters can be inappropriate. For example, Ester et al., (1996) developed the density-based DBSCAN algorithm specifically to handle spatial (GPS) data. Characterized by thin loops and irregular shapes, such data does not fit a center-based clustering paradigm at all. Another example is directional data lying on a sphere (Banerjee et al., (2005), Salah and Nadif, (2019)). In this case, the clusters have centers but are not convex.
@@ 3.5 Generating Archetypes from Natural Language
To fully realize the potential of a high-level, archetype-based approach to synthetic data generation, our package repliclust draws on the OpenAI API (OpenAI, (2023)) to allow users to create data set archetypes directly from verbal descriptions in English.
To map a user’s natural language input to the high-level geometric paramters of Table 1), we use few-shot prompting of OpenAI’s GPT-4o large language model model (Brown et al., (2020), OpenAI et al., (2024)). We use this same approach to automatically create short identifiers for archetypes, such as “ten_highly_oblong_very_different_shapes_moderate_overlap”. These identifiers contain no white space and obey the naming rules for Python variables.
Version 1.0.0. of repliclust uses 22-shot prompting to map archetype descriptions to parameters, and 11-shot prompting to generate archetype identifiers, which appears to work well. In the rare case that the natural language workflow fails, repliclust throws an error. We release all few-shot examples and prompt templates in the natural_language module of our code base. For convenience, they are reproduced in Appendix F.
@@ 4 Results
In this section, we first illustrate the convenience of archetype-based data generation by conducting a mock benchmark between several clustering algorithms. Second, we verify that our notion of clustering overlap effectively captures clustering difficulty (even in higher dimensions), and study the distribution of pairwise overlaps in data sets with multiple clusters.
@@ 4.1 Mock Benchmark
Repliclust maps these descriptions to the following data set archetypes [^3] With one exception: the automatically generated names for archetype 5. and 6. did not end in the suffix “_each.” We added this suffix for clarity.:
@@
    { ‘name’: ‘twelve_clusters_different_distributions’, ‘n_clusters’: 12,
    ‘dim’: 2, ‘n_samples’: 1200, ‘aspect_ref’: 1.5, ‘aspect_maxmin’: 2,
    ‘radius_maxmin’: 3, ‘imbalance_ratio’: 2, ‘max_overlap’: 0.05,
    ‘min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’, ‘gamma’, ‘weibull’, ‘lognormal’] }
@@
    { ‘name’: ‘twelve_different_distributions_high_class_imbalance’,
    ‘n_clusters’: 12, ‘dim’: 2, ‘n_samples’: 1200, ‘aspect_ref’: 1.5,
    ‘aspect_maxmin’: 2, ‘radius_maxmin’: 3, ‘imbalance_ratio’: 5,
    ‘max_overlap’: 0.05, ‘min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’, ‘gamma’, ‘weibull’, ‘lognormal’] }
@@
    { ‘name’: ‘seven_highly_separated_10d_very_different_shapes’,
    ‘n_clusters’: 7, ‘dim’: 10, ‘n_samples’: 700, ‘aspect_ref’: 1.5,
    ‘aspect_maxmin’: 3.0, ‘radius_maxmin’: 3.0, ‘imbalance_ratio’: 2,
    ‘max_overlap’: 0.0001, ‘min_overlap’: 1e-05,
    ‘distributions’: [‘normal’, ‘exponential’] }
@@
    { ‘name’: ‘seven_very_different_shapes_significant_overlap_10d’,
    ‘n_clusters’: 7, ‘dim’: 10, ‘n_samples’: 700, ‘aspect_ref’: 1.5,
    ‘aspect_maxmin’: 3.0, ‘radius_maxmin’: 3, ‘imbalance_ratio’: 2,
    ‘max_overlap’: 0.2, ‘min_overlap’: 0.1,
    ‘distributions’: [‘normal’, ‘exponential’] }
@@
    { ‘name’: ‘four_clusters_100d_100_samples_each’,
    ‘n_clusters’: 4, ‘dim’: 100, ‘n_samples’: 400, ‘aspect_ref’: 1.5,
    ‘aspect_maxmin’: 2, ‘radius_maxmin’: 3, ‘imbalance_ratio’: 2,
    ‘max_overlap’: 0.05, ‘min_overlap’: 0.001,
    ‘distributions’: [‘normal’, ‘exponential’] }
@@
    { ‘name’: ‘four_clusters_100d_1000_samples_each’,
    ‘n_clusters’: 4, ‘dim’: 100, ‘n_samples’: 4000, ‘aspect_ref’: 1.0,
    ‘aspect_maxmin’: 1.0, ‘radius_maxmin’: 3, ‘imbalance_ratio’: 2,
    ‘max_overlap’: 0.05, ‘min_overlap’: 0.001,
    ‘distributions’: [‘normal’, ‘exponential’] }
@@ 4.1.1 Examining the Generated Data Sets
Figure 7 shows four representative data sets with convex clusters drawn from each archetype. Figure 8 shows non-convex clusters resulting from applying the distort function (as described in Section 3.4). For archetypes with dimensionality greater than two, we use t-SNE to visualize the data sets in 2D (van der Maaten and Hinton, (2008)).
@@ 4.1.2 Comparing the Clustering Algorithms
We compare the following clustering algorithms: K-Means, hierarchical (with Ward linkage), spectral (with radial-basis function affinity), HDBSCAN, and expectation maximization for a Gaussian mixture model (EM-GMM). We originally intended to include DBSCAN as well. However, the heuristics we tried for choosing the neighborhood radius (including what is suggested in Sander et al., (1998)) did not work well across the range of dimensionalities we consider, resulting in too many noise points (see also the discussion by Schubert et al., (2017)).
We measure performance in terms of adjusted mutual information (AMI) and adjusted Rand index (ARI), based on the ground truth cluster labels (Vinh et al., (2010), Hubert and Arabie, (1985)). To carry out the benchmark, we sample 10 times from each archetype, resulting in 60 distinct data sets. We repeat this process twice to evaluate performance on convex and non-convex clusters (the latter resulting from applying the distort function on new data sets). Thus, the full benchmark is based on 120 distinct data sets. Table 2 shows the results for K-Means, hierarchical, spectral, and EM-GMM. All these algorithms receive the true number of clusters as an input.
Table 3 separately lists the performance for HDBSCAN. [^4] We used min_samples=5 in the scikit-learn implementation (Pedregosa et al., (2011)). We do this for two reasons. First, as a density-based algorithm, HDBSCAN cannot make use of the true number of clusters, putting it at a disadvantage. Second, HDBSCAN reports a large number of noise points ($\approx$60on average), so that its results are not directly comparable to that of the other algorithms. We report the performance of HDBSCAN based only on the non-noise points. This leads to higher performance numbers, which may compensate for the disadvantage HDBSCAN faces in not knowing the true number of clusters.

### v4

# Natural Language-Based Synthetic Data Generation for Cluster Analysis

@@ Abstract
Cluster analysis relies on effective benchmarks for evaluating and comparing different algorithms. Simulation studies on synthetic data are popular because important features of the data sets, such as the overlap between clusters, or the variation in cluster shapes, can be effectively varied. Unfortunately, creating evaluation scenarios is often laborious, as practitioners must translate higher-level scenario descriptions like “clusters with very different shapes” into lower-level geometric parameters such as cluster centers, covariance matrices, etc. To make benchmarks more convenient and informative, we propose synthetic data generation based on direct specification of high-level scenarios, either through verbal descriptions or high-level geometric parameters. Our open-source Python package repliclust implements this workflow, making it easy to set up interpretable and reproducible benchmarks for cluster analysis. A demo of data generation from verbal inputs is available at demo.repliclust.org.
@@ 1 Introduction
The goal of clustering is to separate data points into groups such that points within a group a more similar to each other than to those outside the group (Cormack,, 1971). In practice, it is often not clear what constitutes a cluster (Hennig,, 2015). As a result, many practitioners evaluate cluster analysis algorithms on synthetic data (Milligan,, 1980; Milligan et al.,, 1983; Tibshirani et al.,, 2001; Steinley and Brusco,, 2008; Steinley and Brusco,, 2011; Van Mechelen et al.,, 2023).
Synthetic data is valuable for two reasons. First, it clearly stipulates which data points belong to which cluster, allowing objective evaluation. Second, it allows independently manipulating different aspects of the data (such as the overlap between clusters or the variability of cluster shapes), which is critical for drawing scientific conclusions about the relative merits of different cluster analysis techniques (Milligan,, 1996).
Unfortunately, setting up benchmarks with synthetic data can be laborious. The process typically involves creating data sets for a number of different scenarios. For example, on benchmarks with convex clusters drawn from probabilistic mixture models, the scenarios may involve “clusters of very different shapes and sizes”, “highly overlapping oblong clusters”, “high-dimensional spherical clusters”, etc. (Tibshirani et al.,, 2001).
Existing data generators do not cater directly to such high-level scenarios. Instead, the user must carefully tune simulation parameters to arrive at the desired scenarios (Steinley and Henson,, 2005; Schubert and Zimek,, 2019; Iglesias et al.,, 2019). While some generators make it easy to control the overlaps between clusters, such high-level control typically does not extend to other aspects like the desired diversity of cluster shapes and sizes (Qiu and Joe,, 2006; Shand et al.,, 2019).
@@ 2 Generating Data from High-Level Archetypes
We implement archetype-based generation by summarizing the overall geometry of probabilistic mixture models in terms of a few high-level parameters, as listed in Table 1. To create an archetype, users can either specify these parameters directly or verbally describe the archetype in English. To map an English description to a precise parameter setting, we use few-shot prompting of a large language model (Brown et al.,, 2020; OpenAI,, 2024).
@@ 3.1 Defining Clusters and Mixture Models
To generate an i.i.d. sample from a cluster, we 1) sample the direction $\hat{\mathbf{x}}$ from the ellipsoid defined by the cluster’s principal axes, and 2) sample the length $||\mathbf{x}||_{2}$ according to the cluster’s univariate distribution $f_{\mathcal{C}}$ (which can be one of many supported distributions including normal, lognormal, exponential, Student’s t, gamma, chi-square, Weibull, Gumbel, F, Pareto, beta, and uniform). To make the spread of each cluster depend only on the lengths of the principal axes, we normalize each univariate distribution so that the $68.2\%$ quantiles of its absolute value is unity. For example, if the univariate distribution is exponential with rate $\lambda$, we would actually sample from a re-scaled random variable $\text{Exp}(\lambda)/q_{0.682}$, where the quantile $q_{0.682}$ satisfies $\mathbb{P}(|\text{Exp}(\lambda)|\leq q_{0.682})=0.682$. This rescaling puts all distributions on the same scale as the multivariate normal distribution, which natively satisfies $\mathbb{P}(|\mathcal{N}(0,1)|\leq 1)\approx 0.682$.
@@ 3.2.1 Measuring Pairwise Overlap
Steinley and Henson, (2005) quantify cluster overlap by computing the full distributional overlap in $p$ dimensions. We prefer our one-dimensional notion in terms of minimax classification error, since high-dimensional Euclidean geometry exhibits a number of counterintuitive phenomena. Specifically, as $p\rightarrow\infty$, the majority of a sphere’s volume becomes concentrated in a thin shell from its surface, and so $p$-dimensional overlap goes to zero unless the marginal overlaps in each dimension approach 100% (Blum et al.,, 2020; Steinley and Henson,, 2005).
Figure 3: Cluster overlap based on the misclassification rate of the best linear classifier, in 1D (left) and 2D (right). The black dashed lines show the decision boundaries corresponding to minimax classification rules between the blue and red clusters, and the gray shaded areas represent classification errors. Cluster overlap $\alpha$ is the total probability mass of the gray areas. Here, $\alpha=14.7\%$ for both the left and right panels.
@@
For a pair of multivariate normal clusters with means $\boldsymbol{\mu}_{1}\neq\boldsymbol{\mu}_{2}$ and the same covariance matrix $\boldsymbol{\Sigma}$, the axis $\boldsymbol{a}_{\text{\tiny{LDA}}}=\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$ minimizes the overall misclassification rate, assuming equal class probabilities (see Hastie et al.,, 2009). Unfortunately, the result does not hold for unequal covariance matrices $\boldsymbol{\Sigma}_{1}\neq\boldsymbol{\Sigma}_{2}$. Thus, we propose the approximation $\boldsymbol{a}_{\text{\tiny{LDA}}}=(\frac{\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}}{2})^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$. Figure 4 verifies that this LDA-based approximation closely matches the exact overlap, as compared with a simpler “center-to-center” (C2C) approximation that uses the difference between the cluster centers as the classification axis (i.e., $\mathbf{a}_{\tiny{C2C}}=\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1}$). In the figure, each data point corresponds to a pair of multivariate normal clusters with the pairwise overlap shown. The full simulation is based on 900 pairs generated from a variety of data set archetypes with different cluster shapes and numbers of dimensions (see code repository for implementation details).
Figure 4: Quality of approximating cluster overlap using our LDA and the simpler center-to-center (C2C) approximations. Both approaches show strong correlations with exact cluster overlap, achieving Pearson correlations $r$ close to 1 (left). However, the C2C method incurs significant relative error, while the LDA approximation typically comes within 10% of the exact overlap (right). The dashed lines indicate estimated conditional means.
@@ 3.4 Making Cluster Shapes More Irregular
In many application domains, ellipsoidal center-based clusters are good models for the data. This is often the case when clusters can be characterized by a “typical” element. For example, in single-cell RNA sequencing, Chen et al., (2020) represent the differences in transcriptional activity across cell types using Gaussian mixture models. In this case, the cluster centers correspond to each cell type’s typical gene expression profile. Similarly, in deep learning, a large language model’s hidden states often form compact clusters that can effectively classify text (Howard and Ruder,, 2018), assess the truthfulness of a response (Azaria and Mitchell,, 2023), and help detect out-of-distribution inputs (Ren et al.,, 2023).
However, convex clusters can be inappropriate. For example, Ester et al., (1996) developed the density-based DBSCAN algorithm specifically to handle spatial (GPS) data. Characterized by thin loops and irregular shapes, such data does not fit a center-based clustering paradigm at all. Another example is directional data lying on a sphere (Banerjee et al.,, 2005; Salah and Nadif,, 2019). In this case, the clusters have centers but are not convex.
@@ 3.5 Generating Archetypes from Natural Language
To fully realize the potential of a high-level, archetype-based approach to synthetic data generation, our package repliclust draws on the OpenAI API (OpenAI,, 2023) to allow users to create data set archetypes directly from verbal descriptions in English.
To map a user’s natural language input to the high-level geometric parameters of Table 1, we use few-shot prompting of OpenAI’s GPT-4o large language model model (Brown et al.,, 2020; OpenAI,, 2024). We use this same approach to automatically create short identifiers for archetypes, such as “ten_highly_oblong_very_different_shapes_moderate_overlap”. These identifiers contain no white space and obey the naming rules for Python variables.
Version 1.0.0. of repliclust uses 22-shot prompting to map archetype descriptions to parameters, and 11-shot prompting to generate archetype identifiers, which appears to work well. In the rare case that the natural language workflow fails, repliclust throws an error. We release all few-shot examples and prompt templates in the natural_language module of our code base. They are also reproduced in Appendix F.
@@ 4 Results
In this section, we first illustrate the convenience of archetype-based data generation by conducting a mock benchmark between several clustering algorithms. Second, we verify that our notion of cluster overlap effectively captures clustering difficulty (even in higher dimensions), and study the distribution of pairwise overlaps in data sets with multiple clusters.
@@ 4.1 Mock Benchmark
Repliclust maps these descriptions to the following data set archetypes:
@@
    {‘name’: ‘twelve_clusters_different_distributions’, n_clusters’: 12, ‘dim’: 2, n_samples’: 1200, ‘aspect_ref’: 1.5, aspect_maxmin’: 2, ‘radius_maxmin’: 3, imbalance_ratio’: 2, ‘max_overlap’: 0.05, min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’, gamma’, ‘weibull’, lognormal’]}’
@@
    {‘name’: ‘twelve_different_distributions_high_class_imbalance’, n_clusters’: 12, ‘dim’: 2, n_samples’: 1200, ‘aspect_ref’: 1.5, aspect_maxmin’: 2, ‘radius_maxmin’: 3, imbalance_ratio’: 5, ‘max_overlap’: 0.05, min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’, gamma’, ‘weibull’, lognormal’]}’
@@
    {‘name’: ‘seven_highly_separated_10d_very_different_shapes’, n_clusters’: 7, ‘dim’: 10, n_samples’: 700, ‘aspect_ref’: 1.5, aspect_maxmin’: 3.0, ‘radius_maxmin’: 3.0, imbalance_ratio’: 2, ‘max_overlap’: 0.0001, min_overlap’: 1e-05, ‘distributions’: [‘normal’, ‘exponential’]}
@@
    {‘name’: ‘seven_very_different_shapes_significant_overlap_10d’, n_clusters’: 7, ‘dim’: 10, n_samples’: 700, ‘aspect_ref’: 1.5, aspect_maxmin’: 3.0, ‘radius_maxmin’: 3, imbalance_ratio’: 2, ‘max_overlap’: 0.2, min_overlap’: 0.1, ‘distributions’: [‘normal’, ‘exponential’]}
@@
    {‘name’: ‘four_clusters_100d_100_samples_each’, n_clusters’: 4, ‘dim’: 100, n_samples’: 400, ‘aspect_ref’: 1.5, aspect_maxmin’: 2, ‘radius_maxmin’: 3, imbalance_ratio’: 2, ‘max_overlap’: 0.05, min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’]}
@@
    {‘name’: ‘four_clusters_100d_1000_samples_each’, n_clusters’: 4, ‘dim’: 100, n_samples’: 4000, ‘aspect_ref’: 1.0, aspect_maxmin’: 1.0, ‘radius_maxmin’: 3, imbalance_ratio’: 2, ‘max_overlap’: 0.05, min_overlap’: 0.001, ‘distributions’: [‘normal’, ‘exponential’]}
Note that the automatically generated names for archetype 5. and 6. did not end in the suffix “_each.” We added this suffix for clarity.
@@ 4.1.1 Examining the Generated Data Sets
Figure 7 shows four representative data sets with convex clusters drawn from each archetype. Figure 8 shows non-convex clusters resulting from applying the distort function (as described in Section 3.4). For archetypes with dimensionality greater than two, we use t-SNE to visualize the data sets in 2D (van der Maaten and Hinton,, 2008).
@@ 4.1.2 Comparing the Clustering Algorithms
We compare the following clustering algorithms: K-Means, hierarchical (with Ward linkage), spectral (with radial-basis function affinity), HDBSCAN, and expectation maximization for a Gaussian mixture model (EM-GMM). We originally intended to include DBSCAN as well. However, the heuristics we tried for choosing the neighborhood radius (including what is suggested in Sander et al.,, 1998) did not work well across the range of dimensionalities we consider, resulting in too many noise points (see also the discussion by Schubert et al.,, 2017).
We measure performance in terms of adjusted mutual information (AMI) and adjusted Rand index (ARI), based on the ground truth cluster labels (Vinh et al.,, 2010; Hubert and Arabie,, 1985). To carry out the benchmark, we sample 10 times from each archetype, resulting in 60 distinct data sets. We repeat this process twice to evaluate performance on convex and non-convex clusters (the latter resulting from applying the distort function on new data sets). Thus, the full benchmark is based on 120 distinct data sets. Table 2 shows the results for K-Means, hierarchical, spectral, and EM-GMM. All these algorithms receive the true number of clusters as an input.
Table 3 separately lists the performance for HDBSCAN, which we ran using the scikit-learn implementation (Pedregosa et al.,, 2011) with min_samples=5. We present the HDBSCAN results separately for two reasons. First, as a density-based algorithm, HDBSCAN cannot make use of the true number of clusters, putting it at a disadvantage. Second, HDBSCAN reports a large number of noise points ($\approx$60on average), so that its results are not directly comparable to that of the other algorithms. We report the performance of HDBSCAN based only on the non-noise points. This leads to higher performance numbers, which may compensate for the disadvantage HDBSCAN faces in not knowing the true number of clusters.